OpenStudy (anonymous):

What is the perimeter of this shape? A. 76 m B. 88 m C. 60 m D. 82 m

OpenStudy (anonymous):

@jordanjamesbay it is A

OpenStudy (anonymous):

how do u know

OpenStudy (anonymous):

bc u see how the 10m and that empty space at the top are the same

OpenStudy (anonymous):

so that would make that a 10 then add them all up together

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

welcome

geerky42 (geerky42):

It's not A.

OpenStudy (anonymous):

then what is it?

OpenStudy (anonymous):

agree with geerky42

OpenStudy (anonymous):

yea it is

geerky42 (geerky42):

How come empty top side and 10m are same?

geerky42 (geerky42):

18-6 = 12, not 10.

OpenStudy (anonymous):

that exacly what i told him

OpenStudy (anonymous):

the top is 12

OpenStudy (anonymous):

no its 10

geerky42 (geerky42):

Figure is not drawn to scale...

OpenStudy (anonymous):

geerky i keep getting A.72 man

OpenStudy (anonymous):

no its 12 do the math man

OpenStudy (anonymous):

18-6=12

OpenStudy (anonymous):

it is A guys

OpenStudy (anonymous):

18 from the bottom and 6 from the short side subtract

geerky42 (geerky42):

Trick behind this is "arrange" sides So it's \(2(26+18)=\boxed{88}\)